### v1

# A functional shunt in the umbilical cord: the role of coiling in solute and heat transfer

Abstract
The umbilical cord plays a critical role in delivering nutrients and oxygen from the placenta to the fetus through the umbilical vein, while the two umbilical arteries carry deoxygenated blood with waste products back to the placenta. Although solute exchange in the placenta has been extensively studied, exchange within the cord tissue has not been investigated. Here, we explore the hypothesis that the coiled structure of the umbilical cord could strengthen diffusive coupling between the arteries and the vein, resulting in a functional shunt. We calculate the diffusion of solutes, such as oxygen, and heat in the umbilical cord to quantify how this shunt is affected by vascular configuration within the cord. We demonstrate that the shunt is enhanced by coiling and vessel proximity. Furthermore, our model predicts that typical vascular configurations of the human cord tend to minimise shunting, which could otherwise disrupt thermal regulation of the fetus. We also show that the exchange, amplified by coiling, can provide additional oxygen supply to the cord tissue surrounding the umbilical vessels.
I Introduction
During pregnancy, fetal development depends on the umbilical cord for delivering nutrients and oxygen from the placenta to the fetus through the umbilical vein (UV) and transferring waste products back through the umbilical arteries (UA), before they are discarded through the maternal circulatory system. Because the umbilical cord plays such a vital role during pregnancy, disruption to blood flow or abnormality in the structure of the cord could significantly affect the growth of the fetus [1, 2].
The umbilical vessels are characterised by their helical geometry (Fig. 1). The umbilical coiling index (UCI), defined as the ratio of the number of coils to the length of the cord, or the inverse helical pitch (Fig. 1a), is commonly used in clinical practice to quantify the coiled geometry [3]. In normal pregnancies, the mean UCI is approximately 0.17 coils/cm [4]; cords with an UCI below the 10th centile (ca. 0.07 coils/cm) are classified as hypocoiled, while cords with an UCI above the 90th centile (ca. 0.3 coils/cm) are classified as hypercoiled (Fig. 1c). Hypercoiling is associated with an increased incidence of fetal growth restriction (FGR), fetal heart deceleration during delivery, vascular thrombosis and cord stenosis [5]. Hypocoiling is associated with a higher incidence of fetal demise, fetal distress at delivery and chromosomal abnormalities [6].
Figure 1: The vascular structure of the umbilical cord. (a) A schematic of the helical structure of the umbilical cord that connects the fetus to the placenta; the pitch $2\pi/\Omega$ represents the average length of a coil relative to the radius of the cord; $\Omega$ is the dimensionless helicity parameter (proportional to the UCI). (b) A histological image of the cross-section of a healthy human cord (adapted from [7] under CC BY-NC 4.0); the vascular cord configuration is characterised by the angle between the umbilical vein and the first artery ($\theta_{1}$, measured relative to the $x$-axis that passes through the centre of the cord and the centre of the vein), the angle between the two arteries $\Delta\theta$, and the artery–vein separation distances $d_{1},\,d_{2}$ normalised by the cord radius. (c) Schematic illustrations of the hypercoiled and hypocoiled umbilical cords.
Given the association between abnormal coiling and adverse pregnancy outcomes, it is important to explore how features of helical cord structure, such as coiling and vascular configuration (Fig. 1b), may affect umbilical cord function. In the umbilical cord, many solutes come from the placenta, where exchange between fetal and maternal blood takes place. While placental exchange has been widely investigated [8, 9, 10, 11], and a few studies have analysed blood flow in umbilical vessels [12, 13, 14, 15], to the best of our knowledge, solute exchange between the umbilical vessels remains unexplored.
In addition to its role in solute transport, the cord also contributes to fetal thermoregulation, which is immature compared to adult physiology, as the fetus lacks the ability to sweat or shiver. Instead, the fetus relies on heat exchange through the feto-placental circulation to maintain an appropriate temperature. Extreme heat exposure or maternal hyperthermia during pregnancy can overwhelm the fetus’s limited ability to regulate its own temperature, potentially leading to complications such as preterm birth, low birth weight and congenital anomalies [16]. A recent computational study by Kasiteropoulou et al. [17] indicated the importance of umbilical coiling in thermal regulation; however, the overall role of cord structure in feto-maternal heat exchange remains an open question.
Finally, in humans, the cord’s connective tissue, known as Wharton’s jelly (WJ), has no extra vasculature besides the umbilical vessels [18]. In other mammals, several smaller vessels have been found through histological studies [19]. It has been hypothesised that they may act as vasa vasorum, smaller vessels that supply the walls of larger blood vessels. Given the coiled nature of the human umbilical vessels, an enhanced exchange could also affect oxygen availability for vessel walls and cells in the extravascular cord tissue.
The objective of this study is therefore to quantify the diffusive coupling of umbilical vessels in the human cord for different solutes and heat by combining a theoretical model with anatomical data. Our model moves beyond the coiling index by encapsulating the complexity of three-dimensional helical vascular configurations in a computationally efficient and interpretable manner, allowing us to define key geometric determinants of functional shunting in the cord for a wide range of solutes. We further contextualise our results for specific structural features of human umbilical cords, based on ex vivo histology and in vivo ultrasound imaging, and show how typically observed vascular configurations tend to minimise the heat exchange between the cord’s vessels. Our model also predicts that cord helicity can enhance oxygen availability to the cord tissue surrounding the umbilical vessels, particularly in the presence of tissue metabolism. The developed methodology provides a building block for future fetal–placental–maternal models, enabling more direct hypothesis testing of structure–function relationships in obstetric pathophysiology, as well as in other complex exchange organs.
II Methods
Our integrative approach is outlined below, starting with a theoretical model that seeks to capture the dominant features of intervascular solute exchange in a coiled umbilical cord.
II.1 Modelling solute exchange between helical vessels
We model the umbilical cord as a straight circular cylinder that hosts three helical vessels (one vein and two arteries). Each vessel has a circular cross-section in the plane orthogonal to the axis $z$ of the cylinder and a constant pitch $2\pi/\Omega$, defined as the length of one complete helical turn parallel to the $z$-axis of the cord, measured relative to the cord’s radius (see Fig. 1a). The dimensionless helicity parameter $\Omega$ is proportional to the UCI. The vascular configuration in the cross-section of a cord (Fig. 1b) can be characterised by the relative distance $d$ between the umbilical arteries and the vein (scaled by the cord’s radius) and the angles between the vein and each of the two arteries $\theta_{1}$ and $\theta_{2}$ (measured with respect to an axis connecting the centre of the cord to the centre of the vein). For simplicity, we assume that both umbilical arteries are of the same size and are the same distance $d$ from the vein (so that $d_{1}=d_{2}=d$ in Fig. 1b).
We model the transfer of solutes (such as oxygen and heat) in the cord tissue ${\color[rgb]{0,0,0}\mathcal{I}}_{\text{t}}$ by a steady diffusion-uptake process and assume the cross-sectional solute concentration to be approximately uniform in each umbilical vessel. We neglect possible differences in solute concentration between the umbilical arterial blood and the amniotic fluid. The governing equations (see Supplement S1 for more details) in normalised dimensionless variables describing solute transport in the umbilical tissue are given by
$\nabla^{2}C$ $=\alpha\left(C+1/\beta\right)\quad\text{in }\;{\color[rgb]{0,0,0}\mathcal{I}}_{\text{t}}\,,$ (1a)
$\eta\,C+(\mathbf{\hat{n}}\cdot\nabla C)$ $=0\quad\text{on }\;{\color[rgb]{0,0,0}\mathcal{I}}_{\text{c}}\,,$ (1b)
$\mathbf{\hat{n}}\cdot\nabla C$ $=0\quad\text{on }\;{\color[rgb]{0,0,0}\mathcal{I}}_{\text{f,\,p}}\,,$ (1c)
[…]
We reduce the three-dimensional problem (1) to two dimensions by exploiting the helical coordinate transformation $X=x\cos{\left(\Omega z\right)}+y\sin{\left(\Omega z\right)}$, $Y=-x\sin{\left(\Omega z\right)}+y\cos{\left(\Omega z\right)}$, $Z=z$, where $z$ measures the distance along the cord’s axis (see Supplement S1 for more details). This transformation allows us to study the three-dimensional effects of helicity in the cross-sectional plane of the cord by solving for $C(X,Y)$, the field that is invariant along the length of the cord relative to a rotating coordinate system. The dimensionless solute flux $N$ per unit length of the cord exchanged between the UV and the two UAs, in the absence of tissue metabolism ($\alpha=0$), and the solute flux per unit length taken up by cord tissue are then given, respectively, by
$N{\color[rgb]{0,0,0}(\Omega,\alpha,\beta,\eta)}=\int_{{\color[rgb]{0,0,0}\partial\mathcal{I}}_{\text{v}}{\color[rgb]{0,0,0}\,\cap\,\mathcal{P}}}\mathbf{\hat{n}}\cdot\left(\nabla_{\perp}C+\Omega^{2}\mathbf{H}\right)\mathop{}\mathrm{d}s\quad\;\text{and}\;\quad N_{\text{u}}=\int_{{\color[rgb]{0,0,0}\mathcal{I}}_{\text{t}}{\color[rgb]{0,0,0}\,\cap\,\mathcal{P}}}\!\left(C+1/\beta\right)\!\mathop{}\mathrm{d}A\,,$ (2)
[…]
$\mathrm{Da}=\frac{D_{\text{t}}L\,N(\Omega,\alpha,\beta,\eta)}{B\,Q}\,,$ (3)
where $D_{\text{t}}$ is the solute diffusivity in tissue, $L$ is the cord length, $Q$ is the total umbilical flow rate and $B$ is an advection-facilitation parameter that accounts for haemoglobin binding of certain solutes, such as oxygen; see Supplement S8 for more details and estimates of $\mathrm{Da}$ for oxygen and heat exchange.
As the metabolic rate $\alpha$ increases, the region surrounding the vessels’ lumens exhibits thin boundary layers and steep concentration gradients. We exclude the effects of these highly localised boundary layers by calculating the extravascular uptake flux within the cord tissue outside a thin strip of thickness $h$ around each vessel (based on typical values for the thickness of the umbilical vessel walls [20]; see Tables S1, S2 in the Supplement). The solute distribution and the corresponding integral fluxes (2) are computed via a finite-element solver of COMSOL ${}^{\text{\texttrademark}}$ 6.0. We validated the use of the 2D field $C(X,Y)$ against direct computationally-expensive solutions of the 3D problem (1) (see Supplement S5 for more details).
II.2 Image analysis of histology and ultrasound data
Histology and ultrasound images of human cord’s cross-sections were taken from published literature [7, 21, 22, 23, 18, 24, 25, 26]. The histology data (stained with hematoxylin and eosin; see Fig. 1b) corresponds to normal pregnancies, with gestational age varying from 37 to 42 weeks; ultrasound data included cords from the third trimester (see Supplementary Table S4 for more details on data sources).
A semi-automated Python script (see Supplement S6) was used to extract structural metrics from the images, including effective radii and distances between each umbilical vessel, as well as the two characteristic angles (Fig. 1b). All measurements were normalised by the cord’s radius such that the scaled cord’s radius for each measurement is of unit length. The shape irregularity of each cord was quantified by the circularity parameter $4\pi\mathcal{A}/(\mathcal{P})^{2}$, where $\mathcal{A}$ is the cross-sectional area and $\mathcal{P}$ is the perimeter of the cord.
III Results
III.1 Cord coiling and vascular proximity promote solute and heat shunting
Figure 2: The impact of cord helicity and vascular proximity on solute exchange. (a) Relative exchange flux $N_{\mathrm{rel}}$ per unit length (4) of the umbilical vein vs. cord helicity $\Omega$ for different UA-UV separation distances $d$, with $\alpha=0,\,\eta=0$. For weakly coiled cords, the solute (heat) exchange amplification scales as $\Omega^{2}$ (relative to the uncoiled vessels of the same length). Concentration fields for four different vessel configurations and three-dimensional diagrams show a hypocoiled ($\Omega$=0.75) and hypercoiled ($\Omega$=3.77) cord, at parameters highlighted with squares on the graph. The colour bar shows solute concentration in normalised units (venous concentration is unity and arterial concentration is zero). (b) Computed exchange flux per unit length of the vein $N$ for small separation $d$. In the limit of small $d$, the leading-order flux is $\mathcal{O}(d^{-1/2})$. Here $\eta=0$. (c) Exchange flux $N$ and relative exchange flux $N_{\mathrm{rel}}$ for different values of the exchange parameter $\eta$ in the cord boundary condition (1b) for $d=0.33$. Other geometric parameters in (a-c) were fixed: $R_{v}=0.25$, $r_{v}=0.25$, $r_{a}=0.12$, $\theta_{1}=0.8\pi$, $\theta_{2}=1.2\pi$.
We first explore how the structural parameters affect solute fluxes between the three vessels in a cord. Fig. 2(a) illustrates the difference in exchange flux for helical ($\Omega>0$) and straight ($\Omega=0$) vessel geometries, in terms of the relative exchange flux
$N_{rel}=\frac{N(\Omega{\color[rgb]{0,0,0},\alpha,\beta,\eta})-N(0{\color[rgb]{0,0,0},\alpha,\beta,\eta})}{N(0{\color[rgb]{0,0,0},\alpha,\beta,\eta})},$ (4)
assuming initially that $\alpha=\eta=0$. Small $\Omega$ corresponds to weakly coiled cords, while a large $\Omega$ corresponds to a tightly coiled cord. This is shown in the 3D representations of one complete coil for $\Omega=0.75$ and $\Omega=3.77$ (Fig. 2a, insets). For large $\Omega$, $N_{rel}$ exceeds unity, implying that the flux can be twice as large as the flux for straight vessels, because tightly coiled cords provide a larger internal surface area for exchange to take place. The concentration field for $\Omega=3.77$, $d=0.3$ (Fig. 2a, inset, top right) varies weakly around the cord perimeter, showing how coiling promotes direct solute exchange between vessels. In contrast, the flux for weakly coiled vessels is similar to the flux for straight vessels. Fig. 2(a) shows that the first correction to the straight vessel flux is of order $\Omega^{2}$. A weak helicity $\Omega\ll 1$ approximation for flux can therefore be obtained by perturbing the three-dimensional problem (1a) in powers of $\Omega^{{\color[rgb]{0,0,0}2}}$.
Since in the problem for $C(X,Y)$ the helicity term appears as $\Omega^{2}$, there is no evidence of chirality in the concentration fields in Fig. 2(a). In the Supplement (Fig. S7b,c), we confirm that two configurations with opposite chirality have the same cross-sectional concentration fields. The position of the arteries in the configuration considered in Fig. 2(a) is symmetrical relative to the $x$-axis. This leads to the symmetry $C(X,Y)=C(X,-Y)$. The Supplement (Fig. S7a) shows concentration fields with increasing helicity for asymmetric configurations. Breaking the symmetry introduces distortion in the concentration fields, but the trend in relative flux is the same as in Fig. 2(a).
Another geometric parameter that contributes to the exchange of solutes is the separation distance $d$ between vessels. The flux is larger for the umbilical vessels that are closer together because concentration gradients are steeper, as shown in Fig. 2(a), colour map inserts. Indeed, Fig. 2(b) illustrates how, as $d$ gets smaller, the flux in the gap between the vessels of $\mathcal{O}(d^{-1/2})$ dominates over the flux outside the gap (see Supplement S3 for more details), and the weak helicity correction of $\mathcal{O}(d{\color[rgb]{0,0,0}{}^{1/2}}\,\Omega^{2})$ provides a good approximation ($d=0.04$; Fig. 2a). On the other hand, for larger $d$, the effect of helicity becomes more dominant, as demonstrated by the larger excess flux $N_{rel}$ (4) relative to the flux in the uncoiled geometry.
Fig. 2(c) illustrates the effects of changing the amniotic exchange parameter $\eta$, which controls the outer boundary condition (1b), on the exchange flux. When $\eta\to 0$, the outer boundary is completely isolated, and the concentration gradients in the cord tissue are driven entirely by the interaction between the umbilical vessels, as in Fig. 2(a,b). When $\eta\to\infty$, the flux increases since the additional sink, the amniotic fluid, forces steeper concentration gradients within the cord tissue. For weakly coiled cords, $N$ almost doubles as $\eta\to\infty$. In contrast, the relative difference in exchange flux $N_{\mathrm{rel}}$ for different $\Omega$ shows that the role of helicity is strongest for $\eta=0$ ($N_{\mathrm{rel}}(\Omega=2.5)-N_{\mathrm{rel}}(\Omega=0.5)\approx 0.6$), compared to $\eta\to\infty$, when this difference falls ($N_{\mathrm{rel}}(\Omega=2.5)-N_{\mathrm{rel}}(\Omega=0.5)\approx 0.2$). We use the no-flux condition ($\eta=0$) for the remainder of this work because the effect of helicity is maximised, providing an upper bound for the relative diffusive coupling $N_{rel}$ between the umbilical vessels in this limit.
III.2 Configurations yielding maximum and minimum exchange
Figure 3: Solute exchange flux $N$ for different cord configurations in the ($\theta_{1}$, $\Delta\theta$) parameter space (see Fig. 1b). The geometrical parameters used were $r_{v}=0.25,r_{a}=0.12,d=0.04,\eta=0,\Omega=3.77$ for (a) $R_{v}=0$ (a straight vein, as shown in (c)) and (b) $R_{v}=0.21$. Cord configurations predicted by the computational model (d) ($\theta_{1}=0.9\pi$; $\Delta\theta=0.2\pi$) that minimise and (e) ($\theta_{1}=0.5\pi$; $\Delta\theta=\pi$) maximise solute exchange between the vessels (shown in (b)). Solute concentrations are plotted in normalised units, varying from zero on the arterial surfaces (blue) to one on the surface of the vein (red). (f) Magnification of a region in (b) showing estimated angles obtained from histology and from ultrasound images of human cord cross-sections from healthy pregnancies. The line of symmetry (dashed) has a slope of $-2$ and a zero intercept at $\theta_{1}=\pi$. (g) and (h) are examples of histology images with irregular shapes, circularity $<1$ ([25], reproduced under CC BY 4.0, and [21], reproduced by permission of Taylor & Francis Ltd, tandfonline.com). (i) Ultrasound image of a cord with approximately unit circularity ([24], reproduced by permission of Karger Publishers, © 2009).
We now consider cross-sectional cord configurations with two arteries in different positions by changing the angles $\theta_{1}$ and $\Delta\theta=|\theta_{2}-\theta_{1}|$ (Fig. 1b). We keep the artery-vein distance ($d=0.04$) constant and use $\theta_{1}\in[0,\pi]$, while $\theta_{2}\in[0,2\pi)$. We plot the exchange flux $N$ for all combinations of $\theta_{1}$ and $\theta_{2}$ for tightly coiled ($\Omega=3.77$) configurations with a straight vein and zero offset $R_{v}=0$ (Fig. 3a) and for a coiled vein with an offset from the cord centreline $R_{v}=0.25$ (Fig. 3b). Configurations with overlapping arteries were excluded from the map; the smallest possible difference in angles such that the arteries do not intersect was $\Delta\theta=0.2\pi$.
When the vein is located at the centre of the cord ($R_{v}=0$; Fig. 3c), the exchange flux is invariant with respect to $\theta_{1}$ and depends only on the relative position $\Delta\theta$ of the two UAs around the UV (Fig. 3a). The exchange is maximised at $\Delta\theta=\pi$ when the arteries are aligned on opposite sides of the vein (Fig. 3c). Increasing the offset of the vein from the cord’s centre ($R_{v}>0$), and hence twisting the vein, amplifies the solute flux (the maximum $N$ increases from about 15 to 26; see Fig. 3b). Furthermore, the displaced vein, which acts as a source, breaks the radial symmetry of the concentration field and produces a complex landscape of enhanced and reduced values for the exchange flux, as shown in Fig. 3(b). The region around the global maximum ($\theta_{1}\approx 0.5\pi$ and $\Delta\theta\approx\pi$) corresponds to the configuration where the arteries are aligned on either side of the vein, perpendicular to the $x$-axis that passes through the centre of the cord and centre of the vein (Fig. 3d). Similarly, local maxima are attained when the arteries are close to each other in the region $\theta_{1}\in[\pi/4,\pi/2]$. The global minimum ($\theta_{1}\approx 0.9\pi$ and $\Delta\theta\approx 0.2\pi$) is achieved when arteries are close to each other and are placed along the $x$-axis near the centre of the cord (Fig. 3e).
The configurations that maximise the exchange flux in both scenarios (Fig. 3c,e) are associated with UAs that coil around the UV, allowing for a larger surface area. However, a more general map in Fig. 3(b) reveals that the exchange flux is smaller when the two UAs are aligned close to the $x$-axis ($\theta_{1}\approx\pi$) than when they are aligned orthogonally to it ($\theta_{1}\approx\pi/2$). This indicates the importance of a specific three-dimensional arrangement of the cord’s vessels, which goes beyond the effective surface area of the UA-UV interface.
To determine realistic configurations of the human umbilical cord, we estimated the angles $\theta_{1}$ and $\Delta\theta$ from histological studies and ultrasound images of umbilical cords from normal pregnancies. Fig. 3(g,h) shows examples of histological images of the human umbilical cord that highlight the irregular, non-circular shape of the cross-section. Shortly after birth, the umbilical arteries undergo constriction in order to prevent fetal blood loss. This effect is seen in the histological images, where the thick arterial walls close off the lumen. Similarly, the lumen of the umbilical vein collapses due to changes in pressure and presents an elliptical cross-section. In contrast, the ultrasound image in Fig. 3(i) shows that an in vivo cord has a more regular circular cross-section. The lumen of the vessels is also more visible and has a circular shape. This supports the model assumption of circular cross-sections for both the cord and the vessels, since it better represents in utero conditions.
Fig. 3(f) shows these measurements plotted in the context of the exchange flux map (Fig. 3b). The observed vascular configurations approximately lie on the line of symmetry $\Delta\theta=-2\theta_{1}+2\pi$ for which the position of the arteries is symmetrical with respect to the $x$-axis, passing through the centre of the cord and the centre of the vein. The model indicates that this approximately symmetric arrangement of the umbilical arteries, with a small $\Delta\theta$, minimises the exchange flux in the 14 human cords from healthy pregnancies that we investigated.
III.3 ‘Virtual’ vasa vasorum
Figure 4: The role of umbilical vascular structure in the cord tissue oxygenation. (a) Uptake flux $N_{u}$ (2) by the cord tissue (excluding a thin strip of thickness $h=0.05$ outside the vascular lumens), for varying rate of tissue metabolism $\alpha$ and cord helicity $\Omega$. The solid lines correspond to $\beta=1$, while the dotted lines show $\beta=2$. Scaled concentration fields for (b) $\Omega=0;\leavevmode\nobreak\ \alpha=1$, (c) $\Omega=4;\leavevmode\nobreak\ \alpha=1$ and (d) $\Omega=4;\leavevmode\nobreak\ \alpha=10^{3}$ plotted in normalised units. (e) The contribution of the umbilical vein to the cord tissue oxygenation, quantified by the uptake flux $N_{s}$ relative to the flux in the case of diffusive oxygen supply by the umbilical arteries alone. Colours indicate different inter-vessel distances $d$. Other geometrical parameters used were: $R_{v}=0.25$, $r_{v}=0.25$, $r_{a}=0.12$, $d=0.3$, $\theta_{1}=0.8\pi$, $\theta_{2}=1.2\pi$ and $\eta=0$.
We hypothesise that the enhanced exchange between helical vessels might boost the supply of oxygen to the walls of the umbilical arteries and hence function as a ‘virtual’ vasa vasorum. In Fig. 4(a) we show the effects of helicity $\Omega$, and the relative solute concentration difference between the UA and the UV, $\beta$ on the extravascular uptake flux $N_{u}$ (see (2)). Similarly to Fig. 2, the uptake flux increases with coiling; however, for weak tissue metabolism, the effect is less noticeable. Fig. 4(b) and (c) show similar effects to Fig. 2(a), where concentration fields become more uniform as $\Omega$ increases. For higher metabolic rate $\alpha$, the impact of helicity gets stronger, and the distortions in the concentration fields in the cord tissue also have a stronger impact on the uptake flux. Fig. 4(d) shows boundary layers around the arteries, which have an elliptical shape due to the effects of helicity on the concentration field. Additionally, the value of $\alpha$ for which the maximum flux is attained at increases for larger $\Omega$.
Fig. 4(a) shows a non-monotonic increase in the extravascular uptake flux as the metabolic rate parameter increases. Clearly, for $\alpha=0$, the uptake flux in the cord tissue must be zero. Large $\alpha$ leads to significant spatial gradients of concentration localised in narrow regions around the vessels (Fig. 4d), reducing solute exchange between the UA and the UV. We assess the impact of metabolism on the oxygen supply to the avascular cord tissue by excluding thin strips of size $h$ (comparable to the vessel wall thickness). Therefore, the uptake flux $N_{u}$ (Fig. 4a) tends to zero for sufficiently strong metabolism ($\alpha\gg 1/h^{2}$), when the tissue oxygenation outside the immediate vicinity of the vascular lumen is negligible. Fig. 4(a) also illustrates the role of the relative venous–arterial concentration difference $\beta$ for the baseline case of $\beta=1$ (oxygen concentration in the UV is twice that of the UA) and for $\beta=2$ (the UV concentration is three times that of the UA). The apparent drop in peak $N_{u}$ at higher $\beta$ masks an increase in the absolute uptake flux that scales with this concentration difference $C_{v}-C_{a}$ (in dimensional variables; see (2)).
To explore the contribution of the umbilical vein to cord tissue oxygenation more directly, we consider a base case in which oxygen in the cord tissue is supplied only by the UA. We compare the net uptake flux $N_{u}$ to this base uptake flux $N_{b}$ by calculating $N_{s}=(N_{u}-N_{b})/N_{b}$, which quantifies the additional oxygen supplied to the tissue due to the presence of the UV (see the Supplement S4 for more details). Fig. 4(e) shows how $N_{s}$ varies with $\alpha$ for different vessel separation $d$. Similarly to Fig. 4(a) there is a non-monotonic increase in excess oxygen supply. As vessels get closer to each other, a large $\alpha$ similarly blocks the interactions between the umbilical vessels and hence diminishes the exchange between them. However, when vessel separation $d$ is small the additional oxygen supply to the cord tissue is twice as large compared to the base case.
IV Discussion
Beyond its obvious role as a conduit for blood enriched with oxygen and nutrients between the placenta and the fetus, rather little is known about the structure–function relationship of the umbilical cord in pregnancy. While functional observations in the cord in vivo are typically limited to ultrasound measurements of blood flow, computational modelling [13, 15] could offer deeper insight into the physiological role of the umbilical cord.
In this study, we examine the hypothesis that there exists a functional shunt between the umbilical vessels through a mathematical model that describes solute diffusion and uptake in the cord tissue. By exploiting a natural transformation of the coordinate system, we encapsulate the complex three-dimensional helical geometry of the umbilical vessels in a computationally efficient and interpretable two-dimensional problem, which is formulated in the cross-sectional plane of the cord. The model reveals that cord coiling and vascular proximity promote solute exchange between the umbilical vessels (Fig. 2). In the limit of moderate helicity ($\Omega\leq 1$), the model shows that the exchange flux is approximately proportional to the square of the cord helicity, or equivalently the UCI augmented by the cord diameter $D$:  $\Omega^{2}\sim{\color[rgb]{0,0,0}D}^{2}(\text{UCI})^{2}$. Similarly, in the limit of a thin gap between the umbilical vessels, separated by the relative distance $d\ll 1$, the model predicts that the dominant contribution to the exchange flux is $\mathcal{O}(d^{-1/2})$. Similar to the earlier results for blood flow modelling in the umbilical cord [13], these geometric determinants show that the UCI alone does not capture the essential features of solute exchange and thus can be a poor indicator of clinical pathology.
The simulations also reveal a non-trivial relationship between the vascular configurations in the cross-section of the umbilical cord and the extent of solute exchange between the vessels (see Fig. 3). We use theoretical predictions to interpret the structural images of cords from healthy pregnancies. In all the samples we investigated the umbilical arteries were positioned approximately symmetrically with respect to the axis, passing through the cord centre and the centre of the vein, and that the observed configurations tend to minimise shunting between the umbilical vessels. To put this in the context of the typical transported solutes, we assess the flux exchanged between the umbilical vessels relative to the flux delivered to the fetus by the umbilical vein, as quantified by a Damköhler number $\mathrm{Da}$ (3) (see Supplement S8). Oxygen transport is diffusion-limited (advection-dominated) in the umbilical cord, (as represented by a small $\mathrm{Da}$), implying that the amount of oxygen exchanged between the cord vessels can be neglected. In contrast, the diffusive coupling for heat transfer may be significant, because the relevant $\mathrm{Da}$ is significantly larger. Given the tendency to minimise this coupling in the measured cord configurations, it opens an intriguing possibility that the cord structure can be partially explained by the demands of fetal thermoregulation.
Although the human umbilical cord tissue lacks vasa vasorum [27], the cord’s WJ contains mesenchymal stem cells (MSCs), fibroblasts, and other connective tissue cells [18], all of which are metabolically active and depend on oxygen to perform their functions, which are essential for the health of the cord (such as tissue repair by the MSCs and collagen synthesis by the fibroblasts). Our model demonstrates how the functional shunt between the umbilical vessels can act as ‘virtual’ vasa vasorum by amplifying the oxygen supply to the cord tissue. While the cells residing in the WJ are likely adapted to a relatively low oxygen environment compared to other tissues [27], we hypothesise that abnormal coiling and vascular proximity could adversely affect the development and function of the cord tissue.
Our modelling approach has many limitations and assumptions that warrant further study. For example, we neglected the effects of potential solute heterogeneity on the surface of umbilical vessels (which is important in parallel uncoiled exchanger systems [28]). However, the length of the cord and the helicity-induced secondary flows [29, 30] will promote the mixing of solutes in umbilical vessels, justifying to some extent the azimuthally uniform approximation. Future work can generalise this model by considering non-zero axial concentration gradients to study exchange along the whole cord and explore more accurately how the functional shunt affects the transport of solutes between the placenta and the fetus.
Additionally, in this study, we primarily analysed the cord structure based on histological images; however, the in vivo structure of the cord is prone to significant perturbations post-delivery. For example, the umbilical vein lumen collapses due to changes in pressure, while the umbilical artery lumen contracts and clots within in help avoid neonatal blood loss [31]. These effects can be seen in Fig. 3(g,h), where the vein has an elliptical shape, the arterial lumens are not readily visible, and the cord’s cross-section can have an irregular (non-circular) shape. In contrast, the ultrasound images better reproduce the in utero environment since the vessels are not collapsed or constricted, with more circular shapes and visible lumens (Fig. 3i). Our analysis relies on the relative centre-of-mass locations of the umbilical vessels, rather than finer shape details, making our results less sensitive to these ex vivo distortions. Nevertheless, in future studies, ultrasound images should be used for a more accurate representation of the umbilical vascular structure.
We have also limited our analysis to the cords from normal human pregnancies. Pathological studies often report thin cords with a significant decrease in WJ [32], and instances of cords with a single UA, which have been associated with fetal growth restriction, pre-eclampsia and other adverse pregnancy outcomes  [33]. Similarly, the morphology of the umbilical cord is known to vary significantly across different placental mammals [19]. These structural differences motivate the need for a systematic quantitative comparison of the cord’s performance in healthy and pathological human pregnancies, as well as its interspecies variations [34, 35].
To conclude, our study offers a new perspective on the structure–function relationship of the umbilical cord by investigating solute exchange between umbilical vessels. Using image-based mathematical modelling, we have demonstrated how cord coiling can significantly enhance diffusive coupling between vessels, while the relative vascular arrangement in human cords typically minimises this coupling. These findings have potential implications for the thermal regulation of the fetus and the oxygenation of the cord tissue. The approach outlined in this paper could be expanded to address the impact of this functional shunt on fetal development in human pregnancy pathologies and across different species.
Ethics. This study did not require ethical approval.
Data Accessibility. All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. Additional data related to this paper may be requested from the authors.
Authors’ Contributions. TW, OEJ and ILC conceptualised the study and co-developed the methodology; EDJ supervised data collection and contributed to the interpretation of the results; SNS supervised image processing and contributed to the interpretation of the results; TW developed the codes, performed the simulations and image analysis, and drafted the initial manuscript. All authors discussed the results and contributed to the writing of the manuscript.
Competing Interests. The authors declare no competing interests.
Funding. This work was supported by the A*STAR Scholarship Award to TW. EDJ, OEJ & ILC also gratefully acknowledge partial support by the Wellcome Leap In Utero Programme.
Acknowledgements. The authors would like to thank Prof. Hwee Kuan Lee and Dr Nicholas Cheng (A*STAR Bioinformatics Institute) for helpful discussions.

### v3

@@
 We also show that the exchange, amplified by coiling, can provide additional oxygen supply to the cord tissue surrounding the umbilical vessels.
+umbilical cord, vascular configuration, helicity, oxygenation, diffusive solute transport, heat exchange
 I Introduction
@@
 (a) A schematic of the helical structure of the umbilical cord that connects the fetus to the placenta; the pitch $2\pi/\Omega$ represents the average length of a coil relative to the radius of the cord; $\Omega$ is the dimensionless helicity parameter (proportional to the UCI).
-(b) A histological image of the cross-section of a healthy human cord (adapted from [7] under CC BY-NC 4.0); the vascular cord configuration is characterised by the angle between the umbilical vein and the first artery ($\theta_{1}$, measured relative to the $x$-axis that passes through the centre of the cord and the centre of the vein), the angle between the two arteries $\Delta\theta$, and the artery–vein separation distances $d_{1},\,d_{2}$ normalised by the cord radius.
-(c) Schematic illustrations of the hypercoiled and hypocoiled umbilical cords.
+(b) A histological image of the cross-section of a healthy human cord (adapted from [7] under $C$ BY-NC 4.0); the vascular cord configuration is characterised by the angle between the umbilical vein and the first artery ($\theta_{1}$, measured relative to the $x$-axis that passes through the centre of the cord and the centre of the vein), the angle between the two arteries $\Delta\theta$, and the artery–vein separation distances $d_{1},\,d_{2}$ normalised by the cord radius.
+(c) Schematic illustrations of hypercoiled and hypocoiled umbilical cords.
 Given the association between abnormal coiling and adverse pregnancy outcomes, it is important to explore how features of helical cord structure, such as coiling and vascular configuration (Fig.
@@
 Our model moves beyond the coiling index by encapsulating the complexity of three-dimensional helical vascular configurations in a computationally efficient and interpretable manner, allowing us to define key geometric determinants of functional shunting in the cord for a wide range of solutes.
-We further contextualise our results for specific structural features of human umbilical cords, based on ex vivo histology and in vivo ultrasound imaging, and show how typically observed vascular configurations tend to minimise the heat exchange between the cord’s vessels.
+We further contextualise our results for specific structural features of human umbilical cords, based on ex vivo histology and in vivo ultrasound imaging, and show how typically observed vascular configurations tend to minimise heat exchange between the cord’s vessels.
 Our model also predicts that cord helicity can enhance oxygen availability to the cord tissue surrounding the umbilical vessels, particularly in the presence of tissue metabolism.
@@
 II.1 Modelling solute exchange between helical vessels
+Parameter Notation Value (cm) Source
+Arterial radius $r^{*}_{a}$ 0.060.20.06-0.2 [20]
+Venous radius $r^{*}_{v}$ 0.10.40.1-0.4 [20]
+Cord radius $R^{*}_{c}$ $0.2-1\,^{\dagger}$ [21]
+Venous helical radius $R^{*}_{v}$ 00.80-0.8 [14]
+Venous wall thickness $h^{*}_{v}$ 0.030.050.03-0.05 [22]
+Arterial wall thickness $h^{*}_{a}$ 0.050.080.05-0.08 [22]
+Cord pitch $2\pi/\Omega^{*}$ 3143-14 [23]
+Cord length $L^{*}$ 5010050-100 [18]
+Artery–vein distance $d^{*}$ 0.040.40.04-0.4 Estimated
+Table 1: Key geometric parameters of the model.
+Stars denote dimensional quantities.
+The pitch was calculated as the inverse umbilical coiling index (UCI) [23], and the artery–vein distance was estimated as described in Supplement S6.
+All the dimensionless lengths in the model are scaled on $R^{*}_{c}=1$ cm.
 We model the umbilical cord as a straight circular cylinder that hosts three helical vessels (one vein and two arteries).
-Each vessel has a circular cross-section in the plane orthogonal to the axis $z$ of the cylinder and a constant pitch $2\pi/\Omega$, defined as the length of one complete helical turn parallel to the $z$-axis of the cord, measured relative to the cord’s radius (see Fig.
+Each vessel has a circular cross-section in the plane orthogonal to the axis $z$ of the cylinder and a constant (and common) pitch $2\pi/\Omega$, defined as the length of one complete helical turn parallel to the $z$-axis of the cord, measured relative to the cord’s radius (see Fig.
 1a).
@@
 1b).
-We model the transfer of solutes (such as oxygen and heat) in the cord tissue ${\color[rgb]{0,0,0}\mathcal{I}}_{\text{t}}$ by a steady diffusion-uptake process and assume the cross-sectional solute concentration to be approximately uniform in each umbilical vessel.
-We neglect possible differences in solute concentration between the umbilical arterial blood and the amniotic fluid.
-The governing equations (see Supplement S1 for more details) in normalised dimensionless variables describing solute transport in the umbilical tissue are given by
-$\nabla^{2}C$ $=\alpha\left(C+1/\beta\right)\quad\text{in }\;{\color[rgb]{0,0,0}\mathcal{I}}_{\text{t}}\,,$ (1a)
-$\eta\,C+(\mathbf{\hat{n}}\cdot\nabla C)$ $=0\quad\text{on }\;{\color[rgb]{0,0,0}\mathcal{I}}_{\text{c}}\,,$ (1b)
-$\mathbf{\hat{n}}\cdot\nabla C$ $=0\quad\text{on }\;{\color[rgb]{0,0,0}\mathcal{I}}_{\text{f,\,p}}\,,$ (1c)
-[…]
-[…]
+We model the transfer of solutes (such as oxygen) and heat in the cord tissue $\mathcal{I}_{\text{t}}$ by a steady diffusion-uptake process and assume the cross-sectional solute concentration to be approximately uniform in each umbilical vessel.
+While blood flow in the umbilical vein is approximately steady, with weak fluctuations [24], arterial flow is pulsatile [12].
+However, the diffusion (with tissue diffusivity $D^{*}_{t}$) of oxygen and heat between the vessels is much slower ($(d^{*})^{2}/D^{*}_{t}\sim 10-10^{3}$ s, for $d^{*}\sim 1$ mm; see Table 1 and Supplement) compared to the typical fetal heart-rate timescale ($10^{-1}-1$ s), justifying the steady-transport approximation.
+We also neglect transmural flow in the cord tissue [25] and possible differences in solute concentration between the umbilical arterial blood and the amniotic fluid.
+The governing equations (see Supplement S1 for more details) in dimensionless variables describing solute transport in the umbilical tissue are given by
+$\nabla^{2}C$ $=\alpha\left(C+1/\beta\right)\quad\text{in }\;\mathcal{I}_{\text{t}}\,$, (1a)
+$\eta\,C+(\mathbf{\hat{n}}\cdot\nabla C)$ $=0\quad\text{on }\;\mathcal{I}_{\text{c}}\,$, (1b)
+$\mathbf{\hat{n}}\cdot\nabla C$ $=0\quad\text{on }\;\mathcal{I}_{\text{f,\,p}}\,$, (1c)
+$C$ $=0\quad\text{on }\;\mathcal{I}_{\text{a${}_{i}$}}\;\text{for }\;i=1,2\,,\quad\text{and}\quad C=1\quad\text{on }\;\mathcal{I}_{\text{v}}\,.$ (1d)
+Here $C(x,y,z)$ is the dimensionless solute concentration (or temperature) normalised by the vein-artery concentration difference $C_{v}-C_{a}$ and centred at the arterial concentration $C_{a}$, so that the dimensional concentration is $C_{a}+(C_{v}-C_{a})C$.
+In (1a), $\alpha$ is a dimensionless metabolic rate parameter (assumed zero for all the solutes but oxygen), and $\beta=(C_{v}-C_{a})/C_{a}$ is the relative solute concentration difference between the umbilical artery (UA) and the umbilical vein (UV) (see Supplement S1).
+Exchange at the interface $\mathcal{I}_{\text{c}}$ between the umbilical cord and the amniotic fluid is characterised by the amniotic exchange parameter $\eta$ in (1b); when $\eta=0$ no solute is lost through outer cord’s surface; $\eta\to\infty$ in (1b) enforces $C=0$ at the cord–amniotic interface.
+We impose no-flux conditions in (1c) at the tissue cross-sections bounding the fetal and placental ends of the umbilical cord respectively ($\mathcal{I}_{\text{f,\,p}}$); here $\mathbf{\hat{n}}$ denotes a unit normal pointing outside the domain occupied by the cord tissue $\mathcal{I}_{\text{t}}$.
+Finally, assuming zero axial concentration gradient along the cord, we set fixed solute concentrations on the surface of the umbilical arteries $\mathcal{I}_{\text{a${}_{1}$}}$, $\mathcal{I}_{\text{a${}_{2}$}}$ and the vein $\mathcal{I}_{\text{v}}$ in (1d), so that we can focus attention on intervascular exchange through cord tissue, accounting for its internal helical structure.
 We reduce the three-dimensional problem (1) to two dimensions by exploiting the helical coordinate transformation $X=x\cos{\left(\Omega z\right)}+y\sin{\left(\Omega z\right)}$, $Y=-x\sin{\left(\Omega z\right)}+y\cos{\left(\Omega z\right)}$, $Z=z$, where $z$ measures the distance along the cord’s axis (see Supplement S1 for more details).
 This transformation allows us to study the three-dimensional effects of helicity in the cross-sectional plane of the cord by solving for $C(X,Y)$, the field that is invariant along the length of the cord relative to a rotating coordinate system.
-The dimensionless solute flux $N$ per unit length of the cord exchanged between the UV and the two UAs, in the absence of tissue metabolism ($\alpha=0$), and the solute flux per unit length taken up by cord tissue are then given, respectively, by
-$N{\color[rgb]{0,0,0}(\Omega,\alpha,\beta,\eta)}=\int_{{\color[rgb]{0,0,0}\partial\mathcal{I}}_{\text{v}}{\color[rgb]{0,0,0}\,\cap\,\mathcal{P}}}\mathbf{\hat{n}}\cdot\left(\nabla_{\perp}C+\Omega^{2}\mathbf{H}\right)\mathop{}\mathrm{d}s\quad\;\text{and}\;\quad N_{\text{u}}=\int_{{\color[rgb]{0,0,0}\mathcal{I}}_{\text{t}}{\color[rgb]{0,0,0}\,\cap\,\mathcal{P}}}\!\left(C+1/\beta\right)\!\mathop{}\mathrm{d}A\,,$ (2)
+We assume that all vessels share the same pitch and that each vessel centreline maintains a fixed distance from the (straight) cord centreline.
+The dimensionless solute flux $N$ per unit length of the cord delivered from the UV to the two UAs and cord tissue, and the solute flux per unit length taken up by cord tissue are then given, respectively, by
+$N(\Omega,\alpha,\beta,\eta)=\int_{\partial\mathcal{I}_{\text{v}}\,\cap\,\mathcal{P}}\mathbf{\hat{n}}\cdot\left(\nabla_{\perp}C+\Omega^{2}\mathbf{H}\right)\mathop{}\mathrm{d}s\quad\;\text{and}\;\quad N_{\text{u}}=\int_{\mathcal{I}_{\text{t}}\,\cap\,\mathcal{P}}\!\alpha\left(C+1/\beta\right)\!\mathop{}\mathrm{d}A\,$, (2)
+where $\nabla_{\perp}\equiv[\partial_{X},\,\partial_{Y},\,0]$ is the 2D gradient operator, and $\mathbf{H}=(Y\,C_{X}-X\,C_{Y})\,[Y,\,-X,\,0]$ is a flux component arising due to cord helicity (see Supplement S2 for further details).
+The integrals in (2) are taken in a plane $\mathcal{P}$, in which $z=\text{constant}$.
+The solute flux $N(\Omega,\alpha,\beta,\eta)$ in transformed coordinates shows explicitly the rotational effects of helicity.
+The helicity factor $\Omega^{2}$ controls the strength of the rotational effects while $\mathbf{H}$ represents how the concentration gradients twist in space.
+This term shows that concentration not only diffuses radially but also experiences angular distortions relative to the vessels.
+We explore below the effects of helicity, vessel proximity and the configuration of the vessels within the cord on the solute fluxes.
+The ratio of a dimensional exchange flux between the umbilical vessels, based on $N$, and the advective flux along the umbilical cord defines a dimensionless Damköhler number
+$\mathrm{Da}=\frac{D^{*}_{t}L^{*}\,N(\Omega,\alpha,\beta,\eta)}{B\,Q^{*}}\,$, (3)
+where $D^{*}_{t}$ is the solute diffusivity in tissue, $L^{*}$ is the cord length, $Q^{*}$ is the total umbilical flow rate and $B$ is an advection-facilitation parameter that accounts for haemoglobin binding of certain solutes, such as oxygen; see Supplement S8 and Table S3 for more details and estimates of $\mathrm{Da}$ for oxygen and heat exchange.
 […]
-$\mathrm{Da}=\frac{D_{\text{t}}L\,N(\Omega,\alpha,\beta,\eta)}{B\,Q}\,,$ (3)
-where $D_{\text{t}}$ is the solute diffusivity in tissue, $L$ is the cord length, $Q$ is the total umbilical flow rate and $B$ is an advection-facilitation parameter that accounts for haemoglobin binding of certain solutes, such as oxygen; see Supplement S8 for more details and estimates of $\mathrm{Da}$ for oxygen and heat exchange.
-As the metabolic rate $\alpha$ increases, the region surrounding the vessels’ lumens exhibits thin boundary layers and steep concentration gradients.
-We exclude the effects of these highly localised boundary layers by calculating the extravascular uptake flux within the cord tissue outside a thin strip of thickness $h$ around each vessel (based on typical values for the thickness of the umbilical vessel walls [20]; see Tables S1, S2 in the Supplement).
-The solute distribution and the corresponding integral fluxes (2) are computed via a finite-element solver of COMSOL ${}^{\text{\texttrademark}}$ 6.0.
-We validated the use of the 2D field $C(X,Y)$ against direct computationally-expensive solutions of the 3D problem (1) (see Supplement S5 for more details).
 II.2 Image analysis of histology and ultrasound data
-Histology and ultrasound images of human cord’s cross-sections were taken from published literature [7, 21, 22, 23, 18, 24, 25, 26].
+Histology and ultrasound images of human cord cross-sections were taken from published literature [7, 26, 27, 28, 18, 29, 30, 31].
 The histology data (stained with hematoxylin and eosin; see Fig.
@@
 (a) Relative exchange flux $N_{\mathrm{rel}}$ per unit length (4) of the umbilical vein vs.
-cord helicity $\Omega$ for different UA-UV separation distances $d$, with $\alpha=0,\,\eta=0$.
+cord helicity $\Omega$ for different UA-UV separation distances $d=0.3$ and $d=0.04$, with $\alpha=0,\,\eta=0$.
 For weakly coiled cords, the solute (heat) exchange amplification scales as $\Omega^{2}$ (relative to the uncoiled vessels of the same length).
-Concentration fields for four different vessel configurations and three-dimensional diagrams show a hypocoiled ($\Omega$=0.75) and hypercoiled ($\Omega$=3.77) cord, at parameters highlighted with squares on the graph.
+Concentration fields for four different vessel configurations and three-dimensional diagrams show a normocoiled ($\Omega$=0.75) and hypercoiled ($\Omega$=3.77) cord, at parameters highlighted with squares on the graph.
+Cords with $\Omega<0.4$ are hypocoiled while cords with $\Omega>2$ are hypercoiled.
 The colour bar shows solute concentration in normalised units (venous concentration is unity and arterial concentration is zero).
-(b) Computed exchange flux per unit length of the vein $N$ for small separation $d$.
+(b) Computed exchange flux per unit length of the vein NN for small separation $d$.
 In the limit of small $d$, the leading-order flux is $\mathcal{O}(d^{-1/2})$.
-Here $\eta=0$.
-(c) Exchange flux $N$ and relative exchange flux $N_{\mathrm{rel}}$ for different values of the exchange parameter $\eta$ in the cord boundary condition (1b) for $d=0.33$.
-Other geometric parameters in (a-c) were fixed: $R_{v}=0.25$, $r_{v}=0.25$, $r_{a}=0.12$, $\theta_{1}=0.8\pi$, $\theta_{2}=1.2\pi$.
+Here $\alpha=0,\,\eta=0$.
+(c) Exchange flux NN and relative exchange flux $N_{\mathrm{rel}}$ for different values of the exchange parameter $\eta$ in the cord boundary condition (1b) for $d=0.33$.
+Other geometric parameters in (a-c) were fixed: $R_{v}=0.25$, $r_{v}=0.25$, $r_{a}=0.12$, $\theta_{1}=0.8\pi$, $\theta_{2}=1.2\pi$ (see Supplement for more details).
+The triangles in (a) and (b) show the relative fold-change in the variables.
 We first explore how the structural parameters affect solute fluxes between the three vessels in a cord.
@@
 2(a) illustrates the difference in exchange flux for helical ($\Omega>0$) and straight ($\Omega=0$) vessel geometries, in terms of the relative exchange flux
-$N_{rel}=\frac{N(\Omega{\color[rgb]{0,0,0},\alpha,\beta,\eta})-N(0{\color[rgb]{0,0,0},\alpha,\beta,\eta})}{N(0{\color[rgb]{0,0,0},\alpha,\beta,\eta})},$ (4)
+$N_{rel}=\frac{N(\Omega,\alpha,\beta,\eta)-N(0,\alpha,\beta,\eta)}{N(0,\alpha,\beta,\eta)},$ (4)
 assuming initially that $\alpha=\eta=0$.
@@
 2(a) shows that the first correction to the straight vessel flux is of order $\Omega^{2}$.
-A weak helicity $\Omega\ll 1$ approximation for flux can therefore be obtained by perturbing the three-dimensional problem (1a) in powers of $\Omega^{{\color[rgb]{0,0,0}2}}$.
+A weak helicity $\Omega\ll 1$ approximation for flux can therefore be obtained by perturbing the three-dimensional problem (1a) in powers of $\Omega^{2}$.
 Since in the problem for $C(X,Y)$ the helicity term appears as $\Omega^{2}$, there is no evidence of chirality in the concentration fields in Fig.
@@
 In the Supplement (Fig.
-S7b,c), we confirm that two configurations with opposite chirality have the same cross-sectional concentration fields.
+S8b,c), we confirm that two configurations with opposite chirality have the same cross-sectional concentration fields.
 The position of the arteries in the configuration considered in Fig.
 2(a) is symmetrical relative to the $x$-axis.
-This leads to the symmetry $C(X,Y)=C(X,-Y)$.
+This leads to the symmetry C(X,Y)=C(X,Y)C(X,Y)=C(X,-Y).
 The Supplement (Fig.
-S7a) shows concentration fields with increasing helicity for asymmetric configurations.
+S8a) shows concentration fields with increasing helicity for asymmetric configurations.
 Breaking the symmetry introduces distortion in the concentration fields, but the trend in relative flux is the same as in Fig.
@@
 Indeed, Fig.
-2(b) illustrates how, as $d$ gets smaller, the flux in the gap between the vessels of $\mathcal{O}(d^{-1/2})$ dominates over the flux outside the gap (see Supplement S3 for more details), and the weak helicity correction of $\mathcal{O}(d{\color[rgb]{0,0,0}{}^{1/2}}\,\Omega^{2})$ provides a good approximation ($d=0.04$; Fig.
+2(b) illustrates how, as $d$ gets smaller, the flux in the gap between the vessels of $\mathcal{O}(d^{-1/2})$ dominates over the flux outside the gap (see Supplement S3 for more details), and the weak helicity correction of $\mathcal{O}(d^{1/2}\,\Omega^{2})$ provides a good approximation ($d=0.04$; Fig.
 2a).
@@
 We use the no-flux condition ($\eta=0$) for the remainder of this work because the effect of helicity is maximised, providing an upper bound for the relative diffusive coupling $N_{rel}$ between the umbilical vessels in this limit.
+For the range of exchange fluxes $N$ predicted in Fig.
+2, the characteristic values of the Damköhler number (3), which quantifies the extent of ‘shunting’ between the UV and UA, are $\mathrm{Da}_{\text{O}_{2}}\sim 10^{-5}$ and $\mathrm{Da}_{\text{heat}}\sim 10^{-1}$ for oxygen and heat, respectively (see Supplement S8 for more details).
+Thus, although no physiological shunting is expected for oxygen exchange, heat exchange could be sensitive to the cord configuration.
 III.2 Configurations yielding maximum and minimum exchange
@@
 1b).
-The geometrical parameters used were $r_{v}=0.25,r_{a}=0.12,d=0.04,\eta=0,\Omega=3.77$ for (a) $R_{v}=0$ (a straight vein, as shown in (c)) and (b) $R_{v}=0.21$.
-Cord configurations predicted by the computational model (d) ($\theta_{1}=0.9\pi$; $\Delta\theta=0.2\pi$) that minimise and (e) ($\theta_{1}=0.5\pi$; $\Delta\theta=\pi$) maximise solute exchange between the vessels (shown in (b)).
-Solute concentrations are plotted in normalised units, varying from zero on the arterial surfaces (blue) to one on the surface of the vein (red).
+The geometrical parameters used were $r_{v}=0.25,r_{a}=0.12,d=0.04,\eta=0,\Omega=3.77$ (see Table 1) for (a) $R_{v}=0$ (a straight vein, as shown in (c)) and (b) $R_{v}=0.21$.
+Cord configurations predicted by the computational model (e) ($\theta_{1}=0.9\pi$; $\Delta\theta=0.2\pi$) that minimise and (d) ($\theta_{1}=0.5\pi$; $\Delta\theta=\pi$) maximise solute exchange between the vessels (shown in (b)).
+Solute concentrations are plotted in normalised units, varying from $C=0$ on the arterial surfaces (blue) to $C=1$ on the surface of the vein (red).
 (f) Magnification of a region in (b) showing estimated angles obtained from histology and from ultrasound images of human cord cross-sections from healthy pregnancies.
-The line of symmetry (dashed) has a slope of $-2$ and a zero intercept at $\theta_{1}=\pi$.
-(g) and (h) are examples of histology images with irregular shapes, circularity $<1$ ([25], reproduced under CC BY 4.0, and [21], reproduced by permission of Taylor & Francis Ltd, tandfonline.com).
+The line of symmetry (dashed) has a slope of 2-2 and a zero intercept at $\theta_{1}=\pi$.
+(g) and (h) are examples of histology images with irregular shapes, circularity <1<1 ([30], reproduced under $C$ BY 4.0, and [26], reproduced by permission of Taylor & Francis Ltd, tandfonline.com).
 (i) Ultrasound image of a cord with approximately unit circularity ([29], reproduced by permission of Karger Publishers, © 2009).
@@
 We plot the exchange flux $N$ for all combinations of $\theta_{1}$ and $\theta_{2}$ for tightly coiled ($\Omega=3.77$) configurations with a straight vein and zero offset $R_{v}=0$ (Fig.
-3a) and for a coiled vein with an offset from the cord centreline $R_{v}=0.25$ (Fig.
-3b).
+3a,c) and for a coiled vein with an offset from the cord centreline $R_{v}=0.25$ (Fig.
+3b,d).
 Configurations with overlapping arteries were excluded from the map; the smallest possible difference in angles such that the arteries do not intersect was $\Delta\theta=0.2\pi$.
@@
 3c).
-Increasing the offset of the vein from the cord’s centre ($R_{v}>0$), and hence twisting the vein, amplifies the solute flux (the maximum $N$ increases from about 15 to 26; see Fig.
+Increasing the offset (helical radius) of the vein from the cord’s centre ($R_{v}>0$), and hence twisting the vein, amplifies the solute flux (the maximum $N$ increases from about 15 to 26; see Fig.
 3b).
@@
 The configurations that maximise the exchange flux in both scenarios (Fig.
-3c,e) are associated with UAs that coil around the UV, allowing for a larger surface area.
+3c,d) are associated with UAs that straddle the UV, allowing for a larger surface area.
 However, a more general map in Fig.
-3(b) reveals that the exchange flux is smaller when the two UAs are aligned close to the $x$-axis ($\theta_{1}\approx\pi$) than when they are aligned orthogonally to it ($\theta_{1}\approx\pi/2$).
+3(b) reveals that the exchange flux is smaller when the two UAs are aligned close to the $x$-axis ($\theta_{1}\approx\pi$;  Fig.
+3e) than when they are aligned orthogonally to it ($\theta_{1}\approx\pi/2$).
 This indicates the importance of a specific three-dimensional arrangement of the cord’s vessels, which goes beyond the effective surface area of the UA-UV interface.
@@
 In Fig.
-4(a) we show the effects of helicity $\Omega$, and the relative solute concentration difference between the UA and the UV, $\beta$ on the extravascular uptake flux $N_{u}$ (see (2)).
+4(a) we show the effects of helicity $\Omega$, and the relative solute concentration difference between the UA and the UV $\beta$, on the extravascular uptake flux $N_{u}$ (see (2)).
 Similarly to Fig.
@@
 4(d) shows boundary layers around the arteries, which have an elliptical shape due to the effects of helicity on the concentration field.
-Additionally, the value of $\alpha$ for which the maximum flux is attained at increases for larger $\Omega$.
+Additionally, the value of $\alpha$ for which the maximum flux is attained increases for larger $\Omega$.
 Fig.
-4(a) shows a non-monotonic increase in the extravascular uptake flux as the metabolic rate parameter increases.
+4(a) shows a non-monotonic increase in the extravascular uptake flux $N_{u}$ (2) as the metabolic rate parameter increases.
 Clearly, for $\alpha=0$, the uptake flux in the cord tissue must be zero.
@@
 Fig.
-4(e) shows how $N_{s}$ varies with $\alpha$ for different vessel separation $d$.
+4(e) shows how $N_{s}$ varies with $\alpha$ for different vessel separations $d$.
 Similarly to Fig.
@@
 As vessels get closer to each other, a large $\alpha$ similarly blocks the interactions between the umbilical vessels and hence diminishes the exchange between them.
-However, when vessel separation $d$ is small the additional oxygen supply to the cord tissue is twice as large compared to the base case.
+However, when the vessel separation $d$ is small the additional oxygen supply to the cord tissue is almost twice as large compared to the base case (see Fig.
+4e).
 IV Discussion
@@
 2).
-In the limit of moderate helicity ($\Omega\leq 1$), the model shows that the exchange flux is approximately proportional to the square of the cord helicity, or equivalently the UCI augmented by the cord diameter $D$:  $\Omega^{2}\sim{\color[rgb]{0,0,0}D}^{2}(\text{UCI})^{2}$.
+In the limit of moderate helicity ($\Omega\leq 1$), the model shows that the exchange flux is approximately proportional to the square of the cord helicity, or equivalently the UCI augmented by the cord diameter $D$:  $\Omega^{2}\sim D^{2}(\text{UCI})^{2}$.
 Similarly, in the limit of a thin gap between the umbilical vessels, separated by the relative distance $d\ll 1$, the model predicts that the dominant contribution to the exchange flux is $\mathcal{O}(d^{-1/2})$.
-Similar to the earlier results for blood flow modelling in the umbilical cord [13], these geometric determinants show that the UCI alone does not capture the essential features of solute exchange and thus can be a poor indicator of clinical pathology.
+Similar to the earlier results for blood flow modelling in the umbilical cord [13, 32], these geometric determinants show that the UCI alone does not capture the essential features of solute exchange and thus can be a poor indicator of clinical pathology.
 The simulations also reveal a non-trivial relationship between the vascular configurations in the cross-section of the umbilical cord and the extent of solute exchange between the vessels (see Fig.
@@
 We use theoretical predictions to interpret the structural images of cords from healthy pregnancies.
-In all the samples we investigated the umbilical arteries were positioned approximately symmetrically with respect to the axis, passing through the cord centre and the centre of the vein, and that the observed configurations tend to minimise shunting between the umbilical vessels.
-To put this in the context of the typical transported solutes, we assess the flux exchanged between the umbilical vessels relative to the flux delivered to the fetus by the umbilical vein, as quantified by a Damköhler number $\mathrm{Da}$ (3) (see Supplement S8).
-Oxygen transport is diffusion-limited (advection-dominated) in the umbilical cord, (as represented by a small $\mathrm{Da}$), implying that the amount of oxygen exchanged between the cord vessels can be neglected.
+In all the samples we investigated, the umbilical arteries were positioned approximately symmetrically with respect to the axis, passing through the cord centre and the centre of the vein, so that the observed configurations tend to minimise shunting between the umbilical vessels.
+To put this in the context of the typical transported solutes, we assessed the flux exchanged between the umbilical vessels relative to the flux delivered to the fetus by the umbilical vein, as quantified by a Damköhler number $\mathrm{Da}$ (see (3)).
+Oxygen transport in the umbilical cord is diffusion-limited (advection-dominated, as represented by a small $\mathrm{Da}$), implying that the amount of oxygen exchanged between the cord vessels can be neglected.
 In contrast, the diffusive coupling for heat transfer may be significant, because the relevant $\mathrm{Da}$ is significantly larger.
@@
 Our model demonstrates how the functional shunt between the umbilical vessels can act as ‘virtual’ vasa vasorum by amplifying the oxygen supply to the cord tissue.
-While the cells residing in the WJ are likely adapted to a relatively low oxygen environment compared to other tissues [27], we hypothesise that abnormal coiling and vascular proximity could adversely affect the development and function of the cord tissue.
+While the cells residing in the WJ are likely adapted to a relatively low oxygen environment compared to other tissues [33], we hypothesise that abnormal coiling and vascular separation distance could adversely affect the development and function of the perivascular cord tissue; however, further experimental studies are needed to evaluate the physiological significance of this mechanism.
 Our modelling approach has many limitations and assumptions that warrant further study.
@@
 However, the length of the cord and the helicity-induced secondary flows [35, 36] will promote the mixing of solutes in umbilical vessels, justifying to some extent the azimuthally uniform approximation.
+We also neglected axial concentration gradients and assumed a constant coiling pitch, whereas human cords can exhibit slow pitch variations [13].
 Future work can generalise this model by considering non-zero axial concentration gradients to study exchange along the whole cord and explore more accurately how the functional shunt affects the transport of solutes between the placenta and the fetus.
 Additionally, in this study, we primarily analysed the cord structure based on histological images; however, the in vivo structure of the cord is prone to significant perturbations post-delivery.
-For example, the umbilical vein lumen collapses due to changes in pressure, while the umbilical artery lumen contracts and clots within in help avoid neonatal blood loss [31].
+For example, the umbilical vein lumen collapses due to changes in pressure, while the umbilical artery lumen contracts and clots within it help avoid neonatal blood loss [37].
 These effects can be seen in Fig.
 3(g,h), where the vein has an elliptical shape, the arterial lumens are not readily visible, and the cord’s cross-section can have an irregular (non-circular) shape.
-In contrast, the ultrasound images better reproduce the in utero environment since the vessels are not collapsed or constricted, with more circular shapes and visible lumens (Fig.
+In contrast, ultrasound images better reproduce the in utero environment since the vessels are not collapsed or constricted, with more circular shapes and visible lumens (Fig.
 3i).
@@
 Nevertheless, in future studies, ultrasound images should be used for a more accurate representation of the umbilical vascular structure.
-We have also limited our analysis to the cords from normal human pregnancies.
+We have also limited our analysis to cords from normal human pregnancies.
 Pathological studies often report thin cords with a significant decrease in WJ [38], and instances of cords with a single UA, which have been associated with fetal growth restriction, pre-eclampsia and other adverse pregnancy outcomes [39].
 Similarly, the morphology of the umbilical cord is known to vary significantly across different placental mammals [19].
-These structural differences motivate the need for a systematic quantitative comparison of the cord’s performance in healthy and pathological human pregnancies, as well as its interspecies variations [34, 35].
+These structural differences motivate the need for a systematic quantitative comparison of the cord’s performance in healthy and pathological human pregnancies (using multiple imaging modalities [40]), as well as its interspecies variations [41, 42].
 To conclude, our study offers a new perspective on the structure–function relationship of the umbilical cord by investigating solute exchange between umbilical vessels.
@@
 All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials.
-Additional data related to this paper may be requested from the authors.
+The associated structural datasets and computational codes can be accessed via the Figshare repository https://doi.org/10.6084/m9.figshare.28382012.
 Authors’ Contributions.
@@
 Competing Interests. The authors declare no competing interests.
-Funding. This work was supported by the A*STAR Scholarship Award to TW. EDJ, OEJ & ILC also gratefully acknowledge partial support by the Wellcome Leap In Utero Programme.
+Funding. This work was supported by an A*STAR Scholarship Award to TW. EDJ, OEJ & ILC also gratefully acknowledge partial support by the Wellcome Leap In Utero Programme.
 Acknowledgements. The authors would like to thank Prof. Hwee Kuan Lee and Dr Nicholas Cheng (A*STAR Bioinformatics Institute) for helpful discussions.